# FMCW Radar Principles and Human Activity Recognition Systems: Foundations, Techniques, and Applications

### v1

@@ 17.1.3 DenseNet for Feature Extraction
5densenet = nn.Sequential(*list(densenet.children())[:-1])
@@ 17.2.5 Fine-tuning Pre-trained Models
6resnet.fc = nn.Linear(resnet.fc.in_features, num_classes)
@@ 17.3.1 ResNet for Radar Image Processing
9resnet.conv1 = nn.Conv2d(1, 64, kernel_size=7, stride=2, padding=3, bias=False)
@@
12resnet = nn.Sequential(*list(resnet.children())[:-1])
@@ 17.3.2 DenseNet for Radar Feature Extraction
5densenet.features.conv0 = nn.Conv2d(1, 64, kernel_size=7, stride=2, padding=3, bias=False)
@@
8densenet = nn.Sequential(densenet.features)
@@ 17.3.3 EfficientNet for Radar Feature Extraction
7efficientnet._conv_stem = nn.Conv2d(1, 32, kernel_size=3, stride=1, padding=1, bias=False)
@@
10efficientnet = nn.Sequential(*list(efficientnet.children())[:-1])
@@ 17.3.5 Fine-tuning Pre-trained Models for Radar Tasks
5efficientnet._conv_stem = nn.Conv2d(1, 32, kernel_size=3, stride=1, padding=1, bias=False)
@@
9efficientnet._fc = nn.Linear(efficientnet._fc.in_features, num_classes)
@@ 17.4.1 Transfer Learning for Radar Tasks
10resnet.fc = nn.Linear(resnet.fc.in_features, num_classes)
@@ 17.4.2 Fine-tuning on Small Radar Datasets
6resnet.fc = nn.Linear(resnet.fc.in_features, num_classes)
@@
Small datasets are prone to overfitting [5], where the model memorizes the training data rather than learning generalizable features. Here are a few strategies to mitigate overfitting when fine-tuning pre-trained models on radar data:
@@ 18.1.1 Key Principles of Contrastive Learning
5class ContrastiveLoss(nn.Module):
@@ 18.2.1 Using CLIP for Radar Feature Extraction
CLIP [34] (Contrastive Language-Image Pre-training) is a contrastive learning model designed to learn multi-modal representations by training on both images and text data. While CLIP was originally developed for vision-language tasks, its contrastive learning framework can be adapted for radar signal processing. By combining radar data with other modalities, such as text descriptions or annotations, CLIP-like models can extract more robust features that capture the relationships between radar signals and additional contextual information.
@@
4class CLIPForRadar(nn.Module):
@@
7 self.radar_encoder = nn.Sequential(
8 nn.Conv2d(1, 64, kernel_size=3, stride=1, padding=1),
9 nn.ReLU(),
10 nn.Conv2d(64, 128, kernel_size=3, stride=1, padding=1),
11 nn.ReLU(),
12 nn.AdaptiveAvgPool2d((1, 1))
@@
14 self.text_encoder = nn.Embedding(1000, text_embed_size) # Example text embedding
15 self.fc_radar = nn.Linear(128, radar_embed_size)
16 self.fc_text = nn.Linear(text_embed_size, radar_embed_size)
@@ 18.2.2 Fine-tuning Contrastive Models on Radar Data
Existing contrastive learning models, such as SimCLR [7] or MoCo [22], can be fine-tuned on radar-specific tasks to improve their performance. These models create a low-dimensional embedding space where similar radar signals are clustered together and different signals are separated. Fine-tuning a pre-trained contrastive learning model on radar data allows the model to adapt its learned representations to the unique characteristics of radar signals, such as noise, clutter, and varying reflection patterns.
@@
SimCLR is a popular contrastive learning method that learns representations by contrasting augmented views of the same image. In the context of radar data, SimCLR can be adapted to compare augmented versions of radar signals [48], such as adding noise, rotating Range-Doppler Maps, or time-shifting signals. Once pre-trained, SimCLR can be fine-tuned on radar data to improve feature extraction for classification and target detection.
1class SimCLR(nn.Module):
@@
4 self.encoder = nn.Sequential(
5 nn.Conv2d(1, 64, kernel_size=3, stride=1, padding=1),
6 nn.ReLU(),
7 nn.Conv2d(64, 128, kernel_size=3, stride=1, padding=1),
8 nn.ReLU(),
9 nn.AdaptiveAvgPool2d((1, 1))
@@
11 self.fc = nn.Linear(128, embed_size)
@@ 19.1 Principles of Traditional Pulse Radar
Traditional pulse radar operates by transmitting short bursts or pulses of electromagnetic energy and measuring the time delay between the transmitted pulse and the returned echo to estimate the distance and velocity of targets [47].
@@ 19.2.1 Resolution Constraints Due to Pulse Width
The resolution of a radar system is determined by the width of the transmitted pulse [60]. In pulse radar systems, shorter pulse widths improve resolution but are technically challenging to generate, especially at long distances. The range resolution ($R_{res}$) of pulse radar is given by:
@@ 20.1 FMCW Radar Principles
FMCW radar [1] operates by transmitting frequency-modulated chirp signals and receiving the reflected signal from a target. The frequency difference (or beat frequency) between the transmitted and received signals encodes information about the target’s distance and velocity.
@@ 20.2 FMCW Signal Processing
The key to FMCW radar’s ability to measure distance and velocity is the signal processing workflow, which involves mixing the transmitted and received signals to obtain the beat frequency. The beat frequency encodes information about the range and velocity of the target. The signal is then processed through an FFT [15] (Fast Fourier Transform) to estimate the target’s range and velocity.
@@ 20.3.1 First FFT: Range Estimation
The first FFT is applied to the beat frequency signal to estimate the range of targets [27]. When the transmitted and received signals are mixed, the resulting beat frequency is proportional to the distance between the radar and the target. By performing an FFT on the time-domain beat signal, the radar can convert it into the frequency domain and identify the peaks that correspond to different target ranges.
@@ 20.3.2 Second FFT: Doppler Estimation
The second FFT is used to estimate the velocity of the target by analyzing phase shifts across multiple radar frames. The Doppler effect [63] causes a frequency shift in the received signal when the target is moving relative to the radar. By comparing the phase of the beat frequency across successive chirps, the second FFT can calculate the Doppler frequency and determine the target’s velocity.
@@ 20.3.3 Third FFT: Range-Doppler Map
The third FFT combines the results of the first and second FFTs to generate a Range-Doppler Map [59], which displays the range and velocity of multiple targets in a two-dimensional plot. The Range-Doppler Map is a powerful tool for multi-target detection and classification, allowing the radar system to visualize both the distance and speed of objects simultaneously.
@@ 21.1 Radar Cube Formation
The radar cube is a 3D matrix [42] that stores information about targets across multiple radar frames. It includes:
@@ 21.1.1 Angle of Arrival (AoA) Estimation
One key component of the radar cube is the angle of arrival [29] (AoA), which provides the direction from which the target’s signal is received. AoA estimation is typically performed using an array of antennas that capture the incoming signal at slightly different times. By measuring the phase difference between signals received at different antennas, the AoA can be determined. Beamforming techniques are commonly used to improve the accuracy of AoA estimation.
@@ 21.2 Point Cloud Generation from Radar Cube
Once the radar cube has been formed, we can use the data within the cube to generate a point cloud. A point cloud represents the spatial distribution of detected targets in 3D space [65]. Each point in the point cloud contains information about:
@@
By converting the range and angle data into Cartesian coordinates, we can visualize the position of the targets in 3D space [46].
@@ 21.2.1 Generating Point Clouds from Radar Cube
The range and angle data from the radar cube can be converted into Cartesian coordinates $(x,y,z)$ using the following formulas [66]:
@@ 21.3 Conclusion
In radar data processing, clustering is an essential technique used to group points that belong to the same target or object in the scene. One of the most effective clustering algorithms for radar point cloud data is DBSCAN [18] (Density-Based Spatial Clustering of Applications with Noise). DBSCAN is particularly useful for radar applications because it can handle noise and does not require the number of clusters to be specified in advance. This chapter introduces the DBSCAN algorithm, explaining how it works and how to apply it to radar data.
@@ 22.1.3 Advantages of DBSCAN for Radar Data
  • No Need to Specify Number of Clusters: Unlike algorithms like k-means [19], DBSCAN does not require the number of clusters to be specified in advance.
@@ 22.3 Applying DBSCAN to Radar Point Cloud Data
In radar data processing, DBSCAN is commonly used to cluster point cloud data, which consists of 3D coordinates representing the spatial distribution of targets. The goal of applying DBSCAN to radar point cloud data is to separate meaningful clusters of points (targets) from noise and outliers [17]. This section will provide a detailed, step-by-step explanation of how to apply DBSCAN to radar point cloud data, including how to tune the key parameters: eps (the distance threshold) and minPts (the minimum number of points to form a cluster).
@@ 22.3.3 Parameter Selection for DBSCAN
Selecting the appropriate values for eps and minPts is critical for achieving optimal clustering performance [8]. Here’s how you can approach parameter selection:
@@ 22.6 Post-Processing and Target Tracking
After applying DBSCAN to cluster radar point cloud data, the next step is post-processing. This phase involves refining the results and tracking the detected objects over time [4]. Post-processing ensures that the clusters produced by DBSCAN are correctly interpreted, noise is minimized, and object trajectories are tracked across multiple frames of radar data. In this section, we will discuss key post-processing techniques and introduce the concept of target tracking, focusing on the integration of Kalman filters to track object movement.
@@ 22.6.2 Introduction to Target Tracking
Target tracking is the process of monitoring the movement of objects [57] (such as vehicles or pedestrians) across multiple frames of radar data. After DBSCAN has identified clusters representing different objects, the next task is to track the movement of these clusters over time. By predicting the future position of an object based on its past trajectory, we can ensure continuous tracking even when the object momentarily leaves the radar’s field of view.
@@ 22.6.3 Tracking Multiple Targets
In many radar applications, it is necessary to track multiple targets simultaneously [67]. This can be achieved by applying a separate Kalman filter to each detected object. The DBSCAN algorithm identifies distinct clusters of points (representing different objects), and each cluster can be tracked using an individual Kalman filter. In cases where objects are lost momentarily (due to occlusion or noise), the Kalman filter can predict the object’s future position, allowing for re-association when the object reappears.
@@ 23.1 Heatmap Generation from Range-Doppler Maps
Range-Doppler Maps [59] are a fundamental output of radar systems, representing the range (distance) and Doppler shift (velocity) of detected targets. These maps provide a two-dimensional representation of the radar scene, where each pixel in the map corresponds to a particular range and velocity bin. By converting Range-Doppler Maps into heatmaps, we can apply deep learning techniques, such as CNNs, to automatically extract features and perform classification or target detection.
@@ 23.1.2 Using CNNs for Target Detection from Heatmaps
Once the Range-Doppler Map is converted into a heatmap, deep learning methods such as Convolutional Neural Networks (CNNs) can be applied to automatically extract spatial features and classify the targets [62]. CNNs are particularly well-suited for this task because they are designed to process two-dimensional data, such as images or heatmaps, by learning patterns and hierarchies of features (e.g., edges, shapes, and textures).
@@
6class CNNModel(nn.Module):
@@
11 self.conv1 = nn.Conv2d(in_channels=1, out_channels=32, kernel_size=3, padding=1) # input: (1, 100, 100)
12 self.pool = nn.MaxPool2d(kernel_size=2, stride=2) # pooling layer to reduce dimensions
@@
15 self.conv2 = nn.Conv2d(32, 64, kernel_size=3, padding=1)
@@
18 self.conv3 = nn.Conv2d(64, 64, kernel_size=3, padding=1)
@@
21 self.fc1 = nn.Linear(64 * 25 * 25, 64) # Assuming input shape (100, 100), reduced to (25, 25) after pooling
22 self.fc2 = nn.Linear(64, num_classes) # 10 target classes
@@
49criterion = nn.CrossEntropyLoss() # For multi-class classification

### v2

@@ 17.1.3 DenseNet for Feature Extraction
5densenet = $n$.Sequential(*list(densenet.children())[:-1])
@@ 17.2.5 Fine-tuning Pre-trained Models
6resnet.fc = $n$.Linear(resnet.fc.in_features, num_classes)
@@ 17.3.1 ResNet for Radar Image Processing
9resnet.conv1 = $n$.Conv2d(1, 64, kernel_size=7, stride=2, padding=3, bias=False)
@@
12resnet = $n$.Sequential(*list(resnet.children())[:-1])
@@ 17.3.2 DenseNet for Radar Feature Extraction
5densenet.features.conv0 = $n$.Conv2d(1, 64, kernel_size=7, stride=2, padding=3, bias=False)
@@
8densenet = $n$.Sequential(densenet.features)
@@ 17.3.3 EfficientNet for Radar Feature Extraction
7efficientnet._conv_stem = $n$.Conv2d(1, 32, kernel_size=3, stride=1, padding=1, bias=False)
@@
10efficientnet = $n$.Sequential(*list(efficientnet.children())[:-1])
@@ 17.3.5 Fine-tuning Pre-trained Models for Radar Tasks
5efficientnet._conv_stem = $n$.Conv2d(1, 32, kernel_size=3, stride=1, padding=1, bias=False)
@@
9efficientnet._fc = $n$.Linear(efficientnet._fc.in_features, num_classes)
@@ 17.4.1 Transfer Learning for Radar Tasks
10resnet.fc = $n$.Linear(resnet.fc.in_features, num_classes)
@@ 17.4.2 Fine-tuning on Small Radar Datasets
6resnet.fc = $n$.Linear(resnet.fc.in_features, num_classes)
@@
Small datasets are prone to overfitting [bornschein2020small], where the model memorizes the training data rather than learning generalizable features. Here are a few strategies to mitigate overfitting when fine-tuning pre-trained models on radar data:
@@ 18.1.1 Key Principles of Contrastive Learning
5class ContrastiveLoss($n$.Module):
@@ 18.2.1 Using CLIP for Radar Feature Extraction
CLIP [ma2024event] (Contrastive Language-Image Pre-training) is a contrastive learning model designed to learn multi-modal representations by training on both images and text data. While CLIP was originally developed for vision-language tasks, its contrastive learning framework can be adapted for radar signal processing. By combining radar data with other modalities, such as text descriptions or annotations, CLIP-like models can extract more robust features that capture the relationships between radar signals and additional contextual information.
@@
4class CLIPForRadar($n$.Module):
@@
7 self.radar_encoder = $n$.Sequential(
8 $n$.Conv2d(1, 64, kernel_size=3, stride=1, padding=1),
9 $n$.ReLU(),
10 $n$.Conv2d(64, 128, kernel_size=3, stride=1, padding=1),
11 $n$.ReLU(),
12 $n$.AdaptiveAvgPool2d((1, 1))
@@
14 self.text_encoder = $n$.Embedding(1000, text_embed_size) # Example text embedding
15 self.fc_radar = $n$.Linear(128, radar_embed_size)
16 self.fc_text = $n$.Linear(text_embed_size, radar_embed_size)
@@ 18.2.2 Fine-tuning Contrastive Models on Radar Data
Existing contrastive learning models, such as SimCLR [chen2020simple] or MoCo [he2020momentum], can be fine-tuned on radar-specific tasks to improve their performance. These models create a low-dimensional embedding space where similar radar signals are clustered together and different signals are separated. Fine-tuning a pre-trained contrastive learning model on radar data allows the model to adapt its learned representations to the unique characteristics of radar signals, such as noise, clutter, and varying reflection patterns.
@@
SimCLR is a popular contrastive learning method that learns representations by contrasting augmented views of the same image. In the context of radar data, SimCLR can be adapted to compare augmented versions of radar signals [rolfsjord2024track], such as adding noise, rotating Range-Doppler Maps, or time-shifting signals. Once pre-trained, SimCLR can be fine-tuned on radar data to improve feature extraction for classification and target detection.
1class SimCLR($n$.Module):
@@
4 self.encoder = $n$.Sequential(
5 $n$.Conv2d(1, 64, kernel_size=3, stride=1, padding=1),
6 $n$.ReLU(),
7 $n$.Conv2d(64, 128, kernel_size=3, stride=1, padding=1),
8 $n$.ReLU(),
9 $n$.AdaptiveAvgPool2d((1, 1))
@@
11 self.fc = $n$.Linear(128, embed_size)
@@ 19.1 Principles of Traditional Pulse Radar
Traditional pulse radar operates by transmitting short bursts or pulses of electromagnetic energy and measuring the time delay between the transmitted pulse and the returned echo to estimate the distance and velocity of targets [richards2010principles].
@@ 19.2.1 Resolution Constraints Due to Pulse Width
The resolution of a radar system is determined by the width of the transmitted pulse [urkowitz1962generalized]. In pulse radar systems, shorter pulse widths improve resolution but are technically challenging to generate, especially at long distances. The range resolution ($R_{res}$) of pulse radar is given by:
@@ 20.1 FMCW Radar Principles
FMCW radar [instrumentsuser] operates by transmitting frequency-modulated chirp signals and receiving the reflected signal from a target. The frequency difference (or beat frequency) between the transmitted and received signals encodes information about the target’s distance and velocity.
@@ 20.2 FMCW Signal Processing
The key to FMCW radar’s ability to measure distance and velocity is the signal processing workflow, which involves mixing the transmitted and received signals to obtain the beat frequency. The beat frequency encodes information about the range and velocity of the target. The signal is then processed through an FFT [duhamel1990fast] (Fast Fourier Transform) to estimate the target’s range and velocity.
@@ 20.3.1 First FFT: Range Estimation
The first FFT is applied to the beat frequency signal to estimate the range of targets [kim2021high]. When the transmitted and received signals are mixed, the resulting beat frequency is proportional to the distance between the radar and the target. By performing an FFT on the time-domain beat signal, the radar can convert it into the frequency domain and identify the peaks that correspond to different target ranges.
@@ 20.3.2 Second FFT: Doppler Estimation
The second FFT is used to estimate the velocity of the target by analyzing phase shifts across multiple radar frames. The Doppler effect [wang2011relativistic] causes a frequency shift in the received signal when the target is moving relative to the radar. By comparing the phase of the beat frequency across successive chirps, the second FFT can calculate the Doppler frequency and determine the target’s velocity.
@@ 20.3.3 Third FFT: Range-Doppler Map
The third FFT combines the results of the first and second FFTs to generate a Range-Doppler Map [tahmoush2011time], which displays the range and velocity of multiple targets in a two-dimensional plot. The Range-Doppler Map is a powerful tool for multi-target detection and classification, allowing the radar system to visualize both the distance and speed of objects simultaneously.
@@ 21.1 Radar Cube Formation
The radar cube is a 3D matrix [palffy2020cnn] that stores information about targets across multiple radar frames. It includes:
@@ 21.1.1 Angle of Arrival (AoA) Estimation
One key component of the radar cube is the angle of arrival [kirkpatrick1953aperture] (AoA), which provides the direction from which the target’s signal is received. AoA estimation is typically performed using an array of antennas that capture the incoming signal at slightly different times. By measuring the phase difference between signals received at different antennas, the AoA can be determined. Beamforming techniques are commonly used to improve the accuracy of AoA estimation.
@@ 21.2 Point Cloud Generation from Radar Cube
Once the radar cube has been formed, we can use the data within the cube to generate a point cloud. A point cloud represents the spatial distribution of detected targets in 3D space [xiang20163d]. Each point in the point cloud contains information about:
@@
By converting the range and angle data into Cartesian coordinates, we can visualize the position of the targets in 3D space [remondino2003point].
@@ 21.2.1 Generating Point Clouds from Radar Cube
The range and angle data from the radar cube can be converted into Cartesian coordinates $(x,y,z)$ using the following formulas [xu2023mmlock]:
@@ 21.3 Conclusion
In radar data processing, clustering is an essential technique used to group points that belong to the same target or object in the scene. One of the most effective clustering algorithms for radar point cloud data is DBSCAN [EsterDen96] (Density-Based Spatial Clustering of Applications with Noise). DBSCAN is particularly useful for radar applications because it can handle noise and does not require the number of clusters to be specified in advance. This chapter introduces the DBSCAN algorithm, explaining how it works and how to apply it to radar data.
@@ 22.1.3 Advantages of DBSCAN for Radar Data
  • No Need to Specify Number of Clusters: Unlike algorithms like k-means [feng2017radar], DBSCAN does not require the number of clusters to be specified in advance.
@@ 22.3 Applying DBSCAN to Radar Point Cloud Data
In radar data processing, DBSCAN is commonly used to cluster point cloud data, which consists of 3D coordinates representing the spatial distribution of targets. The goal of applying DBSCAN to radar point cloud data is to separate meaningful clusters of points (targets) from noise and outliers [ertoz2003finding]. This section will provide a detailed, step-by-step explanation of how to apply DBSCAN to radar point cloud data, including how to tune the key parameters: eps (the distance threshold) and minPts (the minimum number of points to form a cluster).
@@ 22.3.3 Parameter Selection for DBSCAN
Selecting the appropriate values for eps and minPts is critical for achieving optimal clustering performance [cheng2019clustering]. Here’s how you can approach parameter selection:
@@ 22.6 Post-Processing and Target Tracking
After applying DBSCAN to cluster radar point cloud data, the next step is post-processing. This phase involves refining the results and tracking the detected objects over time [bharati2016fast]. Post-processing ensures that the clusters produced by DBSCAN are correctly interpreted, noise is minimized, and object trajectories are tracked across multiple frames of radar data. In this section, we will discuss key post-processing techniques and introduce the concept of target tracking, focusing on the integration of Kalman filters to track object movement.
@@ 22.6.2 Introduction to Target Tracking
Target tracking is the process of monitoring the movement of objects [souza2016target] (such as vehicles or pedestrians) across multiple frames of radar data. After DBSCAN has identified clusters representing different objects, the next task is to track the movement of these clusters over time. By predicting the future position of an object based on its past trajectory, we can ensure continuous tracking even when the object momentarily leaves the radar’s field of view.
@@ 22.6.3 Tracking Multiple Targets
In many radar applications, it is necessary to track multiple targets simultaneously [yu2004collaborative]. This can be achieved by applying a separate Kalman filter to each detected object. The DBSCAN algorithm identifies distinct clusters of points (representing different objects), and each cluster can be tracked using an individual Kalman filter. In cases where objects are lost momentarily (due to occlusion or noise), the Kalman filter can predict the object’s future position, allowing for re-association when the object reappears.
@@ 23.1 Heatmap Generation from Range-Doppler Maps
Range-Doppler Maps [tahmoush2011time] are a fundamental output of radar systems, representing the range (distance) and Doppler shift (velocity) of detected targets. These maps provide a two-dimensional representation of the radar scene, where each pixel in the map corresponds to a particular range and velocity bin. By converting Range-Doppler Maps into heatmaps, we can apply deep learning techniques, such as CNNs, to automatically extract features and perform classification or target detection.
@@ 23.1.2 Using CNNs for Target Detection from Heatmaps
Once the Range-Doppler Map is converted into a heatmap, deep learning methods such as Convolutional Neural Networks (CNNs) can be applied to automatically extract spatial features and classify the targets [wan2019convolutional]. CNNs are particularly well-suited for this task because they are designed to process two-dimensional data, such as images or heatmaps, by learning patterns and hierarchies of features (e.g., edges, shapes, and textures).
@@
6class CNNModel($n$.Module):
@@
11 self.conv1 = $n$.Conv2d(in_channels=1, out_channels=32, kernel_size=3, padding=1) # input: (1, 100, 100)
12 self.pool = $n$.MaxPool2d(kernel_size=2, stride=2) # pooling layer to reduce dimensions
@@
15 self.conv2 = $n$.Conv2d(32, 64, kernel_size=3, padding=1)
@@
18 self.conv3 = $n$.Conv2d(64, 64, kernel_size=3, padding=1)
@@
21 self.fc1 = $n$.Linear(64 * 25 * 25, 64) # Assuming input shape (100, 100), reduced to (25, 25) after pooling
22 self.fc2 = $n$.Linear(64, num_classes) # 10 target classes
@@
49criterion = $n$.CrossEntropyLoss() # For multi-class classification